A Complete Bibliography of the Proceedings Volumes of the ACM Symposia on the Theory of Computing

(1970–1997)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

03 March 2014 Version 1.11

Title word cross-reference

’05 [37]. ’07 [39]. ’08 [40]. ’09 [41].

’13 [45]. 1969 [1]. 1970 [2]. 1971 [3]. 1972 [4]. 1973 [5]. 1974 [6]. 1975 [7]. 1976 [8]. 1977 [9]. 1978 [10]. 1979 [11]. 1980 [12]. 1981 [13]. 1982 [14]. 1983 [15]. 1984 [16]. 1985 [17]. 1986 [18]. 1987 [19]. 1988 [20]. 1989 [21]. 1990 [22]. 1991 [23]. 1992 [24]. 1993 [25]. 1994 [26]. 1995 [27]. 1996 [28]. 1997 [29]. 1998 [30]. 1999 [31].

2001 [33]. 2002 [34]. 2003 [35]. 2004 [36]. 2005 [37]. 2006 [38]. 2007 [39]. 2008 [40]. 2009 [41].

33rd [33]. 36th [36]. 37th [37]. 39th [39].

40th [40].

ACM [35, 37, 38, 39, 40, 41, 42, 43, 44, 45, 7, 8, 9, 10, 12, 13]. Albuquerque [7]. Alto [45]. Angeles [12]. Annual [4, 5, 7, 13, 33, 34, 36, 37, 38, 39, 40, 45, 2, 3, 6, 8, 9, 10, 11, 12, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. April [5, 6, 11, 12, 15, 16]. Atlanta [11, 31]. Austin [5].

Baltimore [22, 37]. Berkeley [18]. Bethesda [41]. Boston [15]. Boulder [9]. British [24, 40].

CA [35, 43]. California [1, 10, 12, 14, 18, 25, 39, 45]. Cambridge [42]. Canada [24, 26, 34, 40]. Chicago [20, 36]. City [19]. Colorado [4, 9]. Columbia [24, 40]. Computing [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. Conference [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Crete [33].

Dallas [30]. DC [16]. del [1]. Denver [4]. Diego [10, 25, 35, 39].

Eighteenth [18]. Eighth [38, 8, 28]. eleventh [11].

fifteenth [15]. Fifth [5, 35, 25, 45]. first [21, 31]. Forty [45]. Forty-fifth [45]. fourteenth [14]. Fourth [4, 34, 24]. Francisco [14].

Georgia [11, 31]. Greece [33].

Heights [3]. Hershey [8]. Hersonissos [33].

Illinois [20, 36]. International [41, 42, 43, 44]. Island [17].

Jose [43]. July [33]. June [27, 35, 36, 39, 41, 42, 43, 45].

Louisiana [23].

MA [42]. Marina [1]. Maryland [22, 37, 41]. Massachusetts [2, 15]. May [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 13, 14, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 34, 37, 38, 40, 41, 44]. Mexico [7]. Milwaukee [13]. Montreal [26, 34].

Nevada [27]. nineteenth [19]. ninth [9, 29]. Northampton [2]. NY [44].

Ohio [3]. Oregon [32]. Orleans [23].

Palo [45]. papers [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Paso [29]. Pennsylvania [8, 28]. Philadelphia [28]. Portland [32]. presented [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Proceedings [38, 42, 43, 44, 45, 12, 13, 37, 39, 40, 41, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Providence [17].

Quebec [26, 34].

record [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Rey [1]. Rhode [17].

San [10, 14, 25, 35, 39, 43]. Seattle [6, 21, 38]. second [2, 22, 32]. seventeenth [17]. Seventh [7, 27]. Shaker [3]. sixteenth [16]. sixth [6, 26]. STOC [37, 39, 40, 41, 45]. STOC’06 [38]. STOC’10 [42]. STOC’11 [43]. STOC’12 [44]. Symposium [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].

tenth [10]. Texas [5, 29, 30]. Theory [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. third [3, 23]. Thirteenth [13]. thirtieth [30]. Thirty [35, 38, 31, 32]. Thirty-Eighth [38]. Thirty-Fifth [35]. thirty-first [31]. Thiry [34]. Thiry-Fourth [34]. twelfth [12]. twentieth [20]. twenty [21, 22, 23, 24, 25, 26, 27, 28, 29]. twenty-eighth [28]. twenty-fifth [25]. twenty-first [21]. twenty-fourth [24]. twenty-ninth [29]. twenty-second [22]. twenty-seventh [27]. twenty-sixth [26].

USA [35, 36, 37, 38, 39, 41, 42, 43, 44, 45].

Vegas [27]. Victoria [24, 40].

WA [38]. Washington [6, 16, 21]. Wisconsin [13].

York [19, 44].

References

ACM:1969:CRA

[1] ACM, editor. Conference record of ACM Symposium on Theory of Computing; papers presented at the symposium, Marina del Rey, California, May 5, 6, 7, 1969. ACM Press, New York, NY, USA, 1969. LCCN QA75.5 .A22 1969.

ACM:1970:CRS

[2] ACM, editor. Conference record of second annual ACM Symposium on Theory of Computing: papers presented at the symposium, Northampton, Massachusetts, May 4, 5, 6, 1970. ACM Press, New York, NY, USA, 1970. LCCN QA75.5 .A22 1970.

ACM:1971:CRT

[3] ACM, editor. Conference record of third annual ACM Symposium on Theory of Computing: papers presented at the symposium, Shaker Heights, Ohio, May 3, 4, 5, 1971. ACM Press, New York, NY, USA, 1971. LCCN QA75.5 .A22 1971.

ACM:1972:CRF

[4] ACM, editor. Conference record, Fourth Annual ACM Symposium on Theory of Computing: papers presented at the symposium, Denver, Colorado, May 1, 2, 3, 1972. ACM Press, New York, NY, USA, 1972. LCCN QA76.6 .A13 1972.

ACM:1973:CRF

[5] ACM, editor. Conference record of Fifth Annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Austin, Texas, April 30–May 2, 1973. ACM Press, New York, NY, USA, 1973. LCCN QA76.6 .A16 1973.

ACM:1974:CRS

[6] ACM, editor. Conference record of sixth annual ACM Symposium on Theory of Computing: papers presented at the symposium, Seattle, Washington, April 30– May 2, 1974. ACM Press, New York, NY, USA, 1974. LCCN QA76.6 .A13 1974.

ACM:1975:CRS

[7] ACM, editor. Conference record of Seventh Annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Albuquerque, New Mexico, May 5–May 7, 1975. ACM Press, New York, NY, USA, 1975. LCCN QA76.6 .A16 1975.

ACM:1976:CRE

[8] ACM, editor. Conference record of the eighth annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Hershey, Pennsylvania, May 3–5, 1976. ACM Press, New York, NY, USA, 1976. LCCN QA 76.6 A12 1976.

ACM:1977:CRN

[9] ACM, editor. Conference record of the ninth annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Boulder, Colorado, May 2–4, 1977. ACM Press, New York, NY, USA, 1977. LCCN QA76.6 .A13 1977.

ACM:1978:CRT

[10] ACM, editor. Conference record of the tenth annual ACM Symposium on Theory of Computing: papers presented at the Symposium, San Diego, California, May 1–3, 1978. ACM Press, New York, NY, USA, 1978. LCCN QA76.6.A13 1978.

ACM:1979:CRE

[11] ACM, editor. Conference record of the eleventh annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Atlanta, Georgia, April 30–May 2, 1979. ACM Press, New York, NY, USA, 1979. LCCN QA75.5.A14 1979.

ACM:1980:CPT

[12] ACM, editor. Conference proceedings of the twelfth annual ACM Symposium on Theory of Computing: papers presented at the symposium, Los Angeles, California, April 28–30, 1980. ACM Press, New York, NY, USA, 1980. ISBN 0-89791-017-6 (paperback). LCCN QA 76.6 A13 1980. US$17.00 (US$14.00 to members). ACM order no. 508800.

ACM:1981:CPT

[13] ACM, editor. Conference proceedings of the Thirteenth Annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Milwaukee, Wisconsin, May 11–13, 1981. ACM Press, New York, NY, USA, 1981. ISBN 0-89791-041-9 (paperback). LCCN QA76.6 .A13 1981. ACM order no. 508810. Also published in Journal of computer and system sciences, vol. 26, no. 3.

ACM:1982:PFA

[14] ACM, editor. Proceedings of the fourteenth annual ACM Symposium on Theory of Computing, San Francisco, California, May 5–7, 1982. ACM Press, New York, NY, USA, 1982. ISBN 0-89791-070-2. LCCN QA75.5 .A14 1982. ACM order no. 508820.

ACM:1983:PFA

[15] ACM, editor. Proceedings of the fifteenth annual ACM Symposium on Theory of Computing, Boston, Massachusetts, April 25–27, 1983. ACM Press, New York, NY, USA, 1983. ISBN 0-89791-099-0. LCCN QA75.5.A14 1983. ACM order no. 508830.

ACM:1984:PSA

[16] ACM, editor. Proceedings of the sixteenth annual ACM Symposium on Theory of Computing, Washington, DC, April 30– May 2, 1984. ACM Press, New York, NY, USA, 1984. ISBN 0-89791-133-4. LCCN QA 76.6 A13 1984. ACM order no. 508840.

ACM:1985:PSA

[17] ACM, editor. Proceedings of the seventeenth annual ACM Symposium on Theory of Computing, Providence, Rhode Island, May 6–8, 1985. ACM Press, New York, NY, USA, 1985. ISBN 0-89791-151-2 (paperback). LCCN QA 76.6 A13 1985. ACM order no. 508850.

ACM:1986:PEA

[18] ACM, editor. Proceedings of the Eighteenth annual ACM Symposium on Theory of Computing, Berkeley, California, May 28–30, 1986. ACM Press, New York, NY, USA, 1986. ISBN 0-89791-193-8. LCCN QA 76.6 A13 1986. ACM order no. 508860.

ACM:1987:PNA

[19] ACM, editor. Proceedings of the nineteenth annual ACM Symposium on Theory of Computing, New York City, May 25–27, 1987. ACM Press, New York, NY, USA, 1987. ISBN 0-89791-221-7 (paperback). LCCN QA 76.6 A13 1987. ACM order no. 508870.

ACM:1988:PTA

[20] ACM, editor. Proceedings of the twentieth annual ACM Symposium on Theory of Computing, Chicago, Illinois, May 2–4, 1988. ACM Press, New York, NY, USA, 1988. ISBN 0-89791-264-0. LCCN QA 76.6 A13 1988. ACM order no. 508880.

ACM:1989:PTF

[21] ACM, editor. Proceedings of the twentyfirst annual ACM Symposium on Theory of Computing, Seattle, Washington, May 15–17, 1989. ACM Press, New York, NY, USA, 1989. ISBN 0-89791-307-8. LCCN QA 76.6 A13 1989. ACM order no. 508890.

ACM:1990:PTS

[22] ACM, editor. Proceedings of the twentysecond annual ACM Symposium on Theory of Computing, Baltimore, Maryland, May 14–16, 1990. ACM Press, New York, NY, USA, 1990. ISBN 0-89791-361-2. LCCN QA76.A15 1990. ACM order no. 508900.

ACM:1991:PTT

[23] ACM, editor. Proceedings of the twenty third annual ACM Symposium on Theory of Computing, New Orleans, Louisiana, May 6–8, 1991. ACM Press, New York, NY, USA, 1991. ISBN 0-89791-397-3. LCCN QA 76.6 A13 1991. ACM order no. 508910.

ACM:1992:PTF

[24] ACM, editor. Proceedings of the twentyfourth annual ACM Symposium on Theory of Computing, Victoria, British Columbia, Canada, May 4–6, 1992. ACM Press, New York, NY, USA, 1992. ISBN 0-89791-511-9. LCCN QA76.A15 1992. ACM order no. 508920.

ACM:1993:PTF

[25] ACM, editor. Proceedings of the twentyfifth annual ACM Symposium on the Theory of Computing, San Diego, California, May 16–18, 1993. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-591-7. LCCN QA 76.6 A13 1993. ACM order no. 508930.

ACM:1994:PTS

[26] ACM, editor. Proceedings of the twentysixth annual ACM Symposium on the Theory of Computing: Montreal, Quebec, Canada, May 23–25, 1994. ACM Press, New York, NY, USA, 1994. ISBN 0-89791-663-8. LCCN QA76 .A15 1994. ACM order no. 508930.

ACM:1995:PTS

[27] ACM, editor. Proceedings of the twentyseventh annual ACM Symposium on Theory of Computing: Las Vegas, Nevada, May 29–June 1, 1995. ACM Press, New York, NY, USA, 1995. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995. ACM order no. 508950.

ACM:1996:PTE

[28] ACM, editor. Proceedings of the twentyeighth annual ACM Symposium on the Theory of Computing, Philadelphia, Pennsylvania, May 22–24, 1996. ACM Press, New York, NY, USA, 1996. ISBN 0-89791-785-5. LCCN QA 76.6 A13 1996.

ACM order no. 508960. Also known as Federated Computing Research Conference (FCRS ’96).

ACM:1997:PTN

[29] ACM, editor. Proceedings of the twentyninth annual ACM Symposium on the Theory of Computing: El Paso, Texas, May 4–6, 1997. ACM Press, New York, NY, USA, 1997. ISBN 0-89791-888-6. LCCN QA76.5 .A849 1997. ACM order no. 508970.

ACM:1998:PTA

[30] ACM, editor. Proceedings of the thirtieth annual ACM Symposium on Theory of Computing: Dallas, Texas, May 23– 26, 1998. ACM Press, New York, NY, USA, 1998. ISBN 0-89791-962-9. LCCN QA75.5 .A14 1998. ACM order number 508980.

ACM:1999:PTF

[31] ACM, editor. Proceedings of the thirtyfirst annual ACM Symposium on Theory of Computing: Atlanta, Georgia, May 1–4, 1999. ACM Press, New York, NY, USA, 1999. ISBN 1-58113-067-8. LCCN QA75.5 .A14 1999. ACM order number 508990.

ACM:2000:PTS

[32] ACM, editor. Proceedings of the thirty second annual ACM Symposium on Theory of Computing: Portland, Oregon, May 21–23, [2000]. ACM Press, New York, NY, USA, 2000. ISBN 1-58113-184-4. ACM order number 508000.

ACM:2001:PAA

[33] ACM, editor. Proceedings of the 33rd Annual ACM Symposium on Theory of Computing: Hersonissos, Crete, Greece, July 6–8, 2001. ACM Press, New York, NY, USA, 2001. ISBN 1-58113-349-9. LCCN QA76.6 .A13 2001. ACM order number 508010.

ACM:2002:PTF

[34] ACM, editor. Proceedings of the Thiry-Fourth Annual ACM Symposium on Theory of Computing, Montréal, Québec, Canada, May 19–21, 2002. ACM Press, New York, NY, USA, 2002. ISBN 1-58113-495-9. LCCN QA75.5 .A22 2002. ACM order number 508020.

ACM:2003:PTF

[35] ACM, editor. Proceedings of the Thirty-Fifth ACM Symposium on Theory of Computing, San Diego, CA, USA, June 9–11, 2003. ACM Press, New York, NY, USA, 2003. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

ACM:2004:PAA

[36] ACM, editor. Proceedings of the 36th Annual ACM Symposium on the Theory of Computing: Chicago, Illinois, USA, June 13–15, 2004. ACM Press, New York, NY, USA, 2004. ISBN 1-58113-852-0. LCCN QA75.5 .A22 2004.

ACM:2005:SPA

[37] ACM, editor. STOC ’05: proceedings of the 37th Annual ACM Symposium on Theory of Computing: Baltimore, Maryland, USA, May 22–24, 2005. ACM Press, New York, NY, USA, 2005. ISBN 1-58113-960-8. LCCN QA75.5 A22 2005.

ACM:2006:SPT

[38] ACM, editor. STOC’06: Proceedings of the Thirty-Eighth Annual ACM Symposium on Theory of Computing 2006, Seattle, WA, USA, May 21–23, 2006. ACM Press, New York, NY, USA, 2006.

ISBN 1-59593-134-1. LCCN QA75.5 .A22

2006. URL http://portal.acm.org/
citation.cfm?id=1132516. ACM order

number 508060.

ACM:2007:SPA

[39] ACM, editor. STOC ’07: proceedings of the 39th Annual ACM Symposium on Theory of Computing, San Diego, California, USA, June 11–13, 2007. ACM Press, New York, NY, USA, 2007. ISBN 1-59593-631-9. LCCN QA75.5 .A22 2007.

ACM:2008:SPA

[40] ACM, editor. STOC ’08: proceedings of the 40th Annual ACM Symposium on Theory of Computing, Victoria, British Columbia, Canada, May 17–20, 2008. ACM Press, New York, NY, USA, 2008. ISBN 1-60558-047-3. LCCN QA76.6 .A152 2008.

ACM:2009:SPA

[41] ACM, editor. STOC ’09: proceedings of the 2009 ACM International Symposium on Theory of Computing, Bethesda, Maryland, USA, May 31–June 2, 2009. ACM Press, New York, NY, USA, 2009. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

ACM:2010:SPA

[42] ACM, editor. STOC’10: Proceedings of the 2010 ACM International Symposium on Theory of Computing: June 5–8, 2010, Cambridge, MA, USA. ACM Press, New York, NY, USA, 2010. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010.

URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

ACM:2011:SPA

[43] ACM, editor. STOC’11: Proceedings of the 2011 ACM International Symposium on Theory of Computing: June 6–8, 2011, San Jose, CA, USA. ACM Press, New York, NY, USA, 2011. ISBN 1-4503-0691-8. LCCN ????

                                        URL http://www.gbv.de/dms/tib-ub-
                                        hannover/63314455x..

ACM:2012:SPA

[44] ACM, editor. STOC’12: Proceedings of the 2012 ACM International Symposium on Theory of Computing: May 19–22, 2012, New York, NY, USA. ACM Press, New York, NY, USA, 2012. ISBN 1-4503-1245-4. LCCN ????

                                        URL http://www.gbv.de/dms/tib-ub-
                                        hannover/63314455x..

ACM:2013:SPF

[45] ACM, editor. STOC ’13: Proceedings of the Forty-fifth Annual ACM Symposium on Theory of Computing: June 1–4, 2013, Palo Alto, California, USA. ACM Press, New York, NY, USA, 2013. ISBN 1-4503-2029-5.